OpenStudy (anonymous):

ILL LOVE U FOREVER Given the following functions f(x) and g(x), solve f[g(7)] and select the correct answer below: f(x) = 4x + 21 g(x) = 2x + 2

OpenStudy (anonymous):

@jamesr @maria12vera

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

85 ?

OpenStudy (anonymous):

First, evaluate g(7). What do you get?

OpenStudy (anonymous):

16 49 85 784 @RunawayGalaxy

OpenStudy (anonymous):

16

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

\[g \left( x \right) = 2x + 2\]\[g \left( 7 \right) = 2\left( 7 \right) + 2\]\[g \left( 7 \right) = ?\]

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Correct, g(7) = 16. Now, the question asks for f(g(7)). So, now you need to calculate f(16). What do you get?

OpenStudy (anonymous):

85

OpenStudy (anonymous):

Correct. Well done!

OpenStudy (anonymous):

THX canu help me with one other one? ill do it in another box to give u 2 medals @ospreytriple

OpenStudy (anonymous):

OK. Go ahead and post.